A chemist added 2L of water to 8L of a 48% acetic acid solution. What is the acid concentration of the new solution?
Answer by dfrazzetto(283)


acetic acid = .48 (8L) = 3.84
New concentration = acid/(new volume) = 3.84 / (10L) = .384 = 38.4%